# Using Diffusion Models to do Data Assimilation

### v1

@@ Abstract
The recent surge in machine learning (ML) methods for geophysical modeling has raised the question of how these same ML methods might be applied to data assimilation (DA). We focus on diffusion modeling (generative artificial intelligence) and on systems that can perform the entire DA, rather than on ML-based tools that are used within an otherwise conventional DA system. We explain that there are (at least) three different types of diffusion-based DA systems and we show in detail that the three systems differ in the type of posterior distribution they target for sampling. The different posterior distributions correspond to different priors and/or likelihoods, which in turn results in different types of training data sets, different computational requirements and different accuracies of their state estimates. We discuss the implications of these findings for the use of diffusion modeling in DA.
@@ 1 Introduction
Data assimilation (DA) is a mathematical and computational framework for updating forecasts in view of observations (see, e.g., Kalnay, 2002). Mathematically, DA relies on Bayes’ rule and all numerical methods for DA can be understood as approximating, in one way or another, a Bayesian posterior distribution. In numerical weather prediction (NWP), we distinguish between Monte Carlo based, ensemble Kalman methods (see, e.g., Evensen, 1994; Anderson, 2001; Tippett et al., 2003; Evensen et al., 2009), optimization-based/variational methods (see, e.g., Talagrand and Courtier, 1987), and “hybrid” methods that combine the Monte Carlo approach with optimization (see, e.g., Hamill and Snyder, 2000; Lorenc, 2003; Zhang et al., 2009; Buehner et al., 2013; Kuhl et al., 2013; Poterjoy and Zhang, 2015). Collectively, we will refer to these methods as “conventional DA,” since these methods have been deployed in NWP for the past few decades with great success.
@@
Recently, there has been an enormous surge in machine learning (ML) methods as applied to geophysical modeling (see, e.g., Price et al., 2025; Lam et al., 2023; Kochkov et al., 2024), which has raised the question of how these same ML tools might be used within or even replace a conventional DA system. The obvious first step might be to replace the physics-based forecast model with a data-driven ML version (see, e.g., Adrian et al., 2025) while continuing to employ conventional DA methods. A more ambitious step is to replace the entire conventional DA system via ML, e.g., using a diffusion model, which is a form of generative artificial intelligence (AI) and will be discussed further below, (see, e.g., Rozet and Louppe, 2023; Chung et al., 2023; Manshausen et al., 2024; Pathak et al., 2024; Qu et al., 2024; Li et al., 2025), through other ML methods (see,e.g., Allen et al., 2025) or even in the absence of a training set (Keller and Potthast, 2024).
In this paper we concentrate on the question: in what ways can a diffusion model replace the entire conventional DA system? We will argue that the answer to this question can be understood by considering subtleties in Bayes’ rule and the different ways of formulating a Bayesian posterior distribution. Briefly, a conventional DA system samples a Bayesian posterior distribution constructed from a prior that is time-dependent and potentially modified by the entire past trajectory of observations. This observation-dependent prior propagates information from previous DA cycles to the current cycle, and we will refer to it as a cycling prior for short. In contrast, we will show that the common practice of using a training set derived from a long time-series of past weather to train diffusion DA algorithm’s samples a different Bayesian posterior distribution with a constant, climatological prior. In some cases, diffusion DA systems attempt to bring in additional information, beyond the observations, in the form of a “forecast,” either generated by the diffusion DA or by other means, but we will show that this is still not equivalent to the posterior distribution obtained using a cycling prior. We are then left to conclude that, except in rare circumstances (Bao et al. (2024)), conventional DA and diffusion DA target different posterior distributions. We feel that a key question that needs to be answered is: which of these posterior distributions is best in the sense that it has the smallest variance? The answer to this question reveals, at least in principle, which algorithmic choices have the best possibility of producing state estimates with the lowest error along with accurate probabilistic inference.
On the other hand, whether a particular DA method is better than another also depends on whether one can accurately sample from that particular posterior distribution in practice. The nonlinearity of the dynamical system being predicted often means that the relevant prior and posterior are non-Gaussian, which implies the potential for significant error in sampling that posterior using conventional DA algorithms that make Gaussian assumptions (Morzfeld and Hodyss (2019)). Hence, in practice the degree of nonlinearity can obscure the differences between the different posterior distributions alluded to above, especially when it is likely that diffusion DA algorithms may be more adept at handling non-Gaussianity than conventional DA. We do not attempt to answer questions about the impact of non-Gaussianity on the differences between these methods because the answer is clearly application dependent. Instead, we focus entirely on explaining the differences between the various possible algorithmic choices in terms of precisely identifying the particular formulation of Bayes’ rule each method is attempting to solve. We emphasize that this identification of which Bayes’ rule each method is attempting to solve will transcend the differences between linear and nonlinear systems as well as any differences between Gaussian and non-Gaussian distributions. Consequently, we will assume that each method is able to accurately sample its own version of Bayes’ rule and the only question left is then about the differences between the various forms of Bayes’ rule. We therefore leave the description of the impact of non-Gaussianity and the application dependent differences between all these methods to future work.
To this end, we focus on a simplified, linear example that is amenable to analysis and analytical expressions (no approximations). Specifically, we show how variants of diffusion DA systems target different Bayesian posterior distributions, defined by different prior distributions and/or likelihoods. We then construct various training sets that imply these distinctly different diffusion models. This process allows us to broadly assess the main differences between various diffusion DA systems and conventional DA.
Our main results are:
  1. 1.
    Traditional diffusion DA is effective at sampling a Bayesian posterior distribution with a fixed, climatological prior, but the accuracy of such a system is inferior to a DA system that samples a Bayesian posterior distribution with a time-evolving cycling prior. Again, whether this result is borne out in practice is likely to strongly depend on the degree of nonlinearity, time between observations, quantity and quality of the observations, etc.
  2. 2.
    A diffusion DA system can sample the exact same posterior distribution as a DA system using a time-evolving cycling prior, but the denoiser in such a diffusion DA system will need to be re-trained at each DA cycle, which incurs a significant computational cost.
  3. 3.
    A diffusion DA system with a fixed, climatological prior can be modified to ingest a forecast in addition to the observations. This forecast appears to add some aspects of the time-evolving cycling prior back into the DA system, but is not entirely equivalent. Furthermore, if this forecast is generated by a separate DA system, the training cost is relatively low and the accuracy is higher than that of a DA system with a climatological prior (but without a forecast). Nevertheless, the resulting accuracy of this extended diffusion DA system that additionally uses a forecast is lower than that of a DA system with a cycling prior.
While these results are rigorous, they directly apply only to a simplified linear system and in the limits of a large training set (for diffusion DA) and a large ensemble size (for conventional DA). Nonetheless, the Bayesian posterior distributions we connect to the different variants of diffusion DA systems generalize to any setup using the same broad algorithmic choices for the diffusion model. We will argue that we can learn a lot from this knowledge of the targeted posterior distributions and that this will allow us to draw practically relevant conclusions.
@@
The rest of this paper is organized as follows. In Section 2 we will introduce both conventional and diffusion-based DA systems. In Section 3 we describe a linear, stochastic dynamical system that will allow us to clearly formulate all aspects of the DA problem analytically. We will apply different forms of diffusion-DA systems to this dynamical model in order to reveal the prior, likelihood and posterior each system corresponds with. In Section 4 we provide a numerical illustration of the main results from Section 3. We close the manuscript with a summary of the major results and their conclusions in Section 5.
2 Conventional and diffusion-based data assimilation
We begin by briefly introducing the fundamental aspects of both conventional DA and diffusion modeling. Our focus here is on revealing how each method samples a different Bayesian posterior distribution.
2.1 Conventional data assimilation
@@
Data assimilation is concerned with approximating a time-evolving Bayesian posterior distribution that describes the probability of a system state $\mathbf{x}_{k}$ at (discrete) time $k$, given observations $\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{k-1},\mathbf{y}_{k}$ up to time $k$:
@@
It is important to note here that information is propagated from cycle-to-cycle by a time evolving prior. In other words, the posterior of the previous cycle is used to generate the prior for the next cycle. For the remainder of this paper we will thus refer to the prior $p(\mathbf{x}_{k}|\mathbf{y}_{1},\dots,\mathbf{y}_{k-1})$ as a cycling prior. Various DA algorithms approximate the above Bayesian posterior distribution in one way or another. Here, we use the ensemble Kalman filter (EnKF) (Burgers et al., 1998; Evensen, 1994; Evensen et al., 2009) as a representative example of a conventional DA system. In preparation for the diffusion modeling to come, we emphasize that the EnKF takes in a training set, referred to in the literature as an “ensemble”, and outputs a new training set that is used at the next cycle.
This regeneration step of the EnKF works as follows. At cycle $k-1$, we have observations $\mathbf{y}_{1},\dots,\mathbf{y}_{k-1}$ and samples from $p(\mathbf{x}_{k-1}|\mathbf{y}_{1},\dots,\mathbf{y}_{k-1})$ in the form of an ensemble $\mathbf{x}_{k-1}^{(i)}$, where superscript $i=1,\dots,n_{e}$ indexes the ensemble members (and $n_{e}$ is the ensemble size). The EnKF makes a forecast for time $k$ by evolving the ensemble $\mathbf{x}_{k-1}^{(i)}$ forward in time using the model; the result is a forecast ensemble, viz.
@@
where $\mathcal{M}(\cdot)$ is a forecast model (physics-based or ML/AI). This forecast ensemble represents the prior, $p(\mathbf{x}_{k}|\mathbf{y}_{1},\dots,\mathbf{y}_{k-1})$, as a collection of samples. The forecast ensemble is updated to an analysis ensemble by employing stochastic ensemble generation (van Leeuwen (2020)), viz.
$\mathbf{x}_{ak}^{(i)}=\mathbf{x}_{fk}^{(i)}+\mathbf{K}(\mathbf{y}_{k+1}-(\mathbf{H}\mathbf{x}_{fk}^{(i)}+\boldsymbol{\varepsilon}^{(i)})),$ (3)
@@
$\mathbf{y}_{k}=\mathbf{H}\mathbf{x}_{k}+\boldsymbol{\varepsilon},$ (4)
where $\boldsymbol{\varepsilon}$ is a Gaussian random variable with mean 00 and covariance matrix $\mathbf{R}$; $\boldsymbol{\varepsilon}^{(i)}$ in (3) is a sample from the same distribution as $\boldsymbol{\varepsilon}$; the matrix
@@
is an ensemble approximation of the Kalman gain and $\mathbf{P}$ is the ensemble covariance (usually localized (see Morzfeld and Hodyss (2023)) and inflated (see Whitaker and Hamill (2012), Hodyss et al. (2016), Gharamti et al. (2019)).
@@ 2.2 Diffusion modeling
The goal in diffusion modeling is to construct a procedure to sample a random variable with probability density function (pdf), $p(\mathbf{x})$, given a sufficiently large number of samples from that distribution in the form of a training set. This training set plays an identical role to the ensemble in the EnKF above. The main distinction between the ensemble in Section 2.2.1 and the training set here is that the training set is never updated with the latest information from observations, but the ensemble of Section 2.2.1 is updated with that information. This distinction will be relaxed in Section 3.3.2.3.2.2 where we develop a diffusion model that does update its training set at each cycle.
The standard approach in diffusion modeling is to set up a forward process in the form of a simple stochastic differential equation (SDE) and to then reverse that process. In the forward process, we start with a sample from $p(\mathbf{x})$ and sequentially add noise to the sample. These steps are then reversed, so that we can obtain a sample from $p(\mathbf{x})$. A neural network is trained on the samples (and the successively noisy versions of it) to enable the reverse process. Once trained, the diffusion model takes in noise and outputs a sample from the desired pdf.
Below we will refer to the state of the forward process with, $\mathbf{u}$, and the state of the reverse process with, $\mathbf{v}$. We employ a very simple forward process following Karras et al. (2022), i.e.,
@@
where the initial conditions for (6) are drawn from $p(\mathbf{x})$ and where $\beta_{t}$ is a standard Brownian motion (i.e. a Wiener process).
This approach is referred to as a “variance-exploding” method because the variance of (6) monotonically increases with time. The solution to (6) has the property that
$\mathbf{u}_{t}|\mathbf{u}_{0}\sim N(\mathbf{u}_{0},t^{2}),$ (7)
which means that the samples from this pdf are created by simply adding Gaussian noise to samples drawn from $p(\mathbf{x})$. This is an extremely important property as we will use (7) to make use of Tweedie’s formula (Efron (2011)) as discussed below.
Given the forward process (6), we have, according to Anderson (1982), a reverse process that is integrated backwards in time (from $t=T$ to $t=0$):
@@
where $p_{t}(\mathbf{v})$ is the time evolution of the marginal pdf of (6) and $\overline{\beta}_{t}$ is a reverse-time Brownian motion. Therefore, we generate the $i^{th}$ sample from $p(\mathbf{x})$ as $\mathbf{x}^{(i)}=\mathbf{v}(t=0)$.
Note that the time, $t$, in both (6) and (8) are not model or physical time. Rather, the parameter, $t$, is to be thought of as, just that, a parameter denoting the virtual time within the diffusion model. In any event, given the highly noisy nature described by (7) we draw the initial condition for (8) from a Gaussian with mean zero and variance $T^{2}$ where $T$ is a large final virtual time of which we imagine the forward process stopped.
The term $\nabla\log p_{t}(\mathbf{v})$ is referred to as the “score function” and we use Tweedie’s formula (see Appendix A) to compute this as
@@
Recall that the conditional expectation $E_{\mathbf{x}\sim p(\mathbf{x}|\mathbf{v}_{t})}[\mathbf{x}]$ is the best estimate in the sense of mean-square of $\mathbf{x}$ for a given $\mathbf{v}$. We can thus calculate the expected value required in (9) by a “denoiser”, $D(\mathbf{v}_{t},t)$, that minimizes a loss function defined from the expected mean squared error, i.e.,
$\ell=E_{\mathbf{x}\sim p(\mathbf{x})}E_{n\sim\mathcal{N}(0,t^{2})}||\mathbf{x}-\mathbf{D}(\mathbf{v}=\mathbf{x}+n,t)||_{2}^{2}$ (11)
@@
where $||\cdot||_{2}^{2}$ denotes the $L_{2}$ norm. In practice, the denoiser is a neural network that is trained to predict $\mathbf{x}$ from $\mathbf{v}_{t}$. The training set is as follows. For each sample $\mathbf{x}$, we have noisy versions at various times in the forward process, which we can obtain by simply adding noise to the sample $\mathbf{x}$ according to the rule defined by (7). These “sample” and “noisy sample” pairs are used to train the denoiser. Once trained, we can use the denoiser to simulate the reverse process:
$\text{d}\mathbf{v}=-2\frac{D(\mathbf{v},t)-\mathbf{v}}{t}\text{d}t+\sqrt{2t}\text{d}\overline{\beta}_{t}.$ (12)
The reverse process now allows us to sample the desired pdf, $p(\mathbf{x})$, by initializing the reverse process with white noise and simulating it backwards in virtual time.
2.3 Diffusion-based data assimilation
Diffusion modeling, as outlined just above, is an ML technique that samples a random variable by fitting a stochastic differential equation (SDE) to a training set (Sohl-Dickstein et al. (2015), Ho et al. (2020)). There is a very large literature on diffusion modeling. A common extension of the description of diffusion modeling given above is conditional image generation in which one generates images of a requested scene given text prompts (see, e.g., Zhan et al., 2024; Ding et al., 2025).

### v2

@@ Abstract
The recent surge in machine learning (ML) methods for geophysical modeling has raised the question of how these methods might be applied to data assimilation (DA). We focus on diffusion modeling (a form of generative artificial intelligence) for systems that can perform the entire DA process, rather than on ML-based tools used within a conventional DA system. We identify at least three distinct types of diffusion-based DA systems and show that they differ in the posterior distribution they target for sampling. These posterior distributions correspond to different priors and/or likelihoods, which in turn result in unique training datasets, computational requirements, and state estimate qualities. Our analysis further shows that a diffusion DA system designed to target the same posterior distribution as current ensemble DA algorithms requires re-training at each DA cycle, which is computationally costly. We discuss the implications of these findings for the use of diffusion modeling in DA.
@@ 1 Introduction
Data assimilation (DA) is a mathematical and computational framework for updating forecasts in view of observations (see, e.g., Kalnay, 2002). Mathematically, DA relies on Bayes’ rule and all numerical methods for DA can be understood as approximating, in one way or another, a Bayesian posterior distribution. In numerical weather prediction (NWP), we distinguish between ensemble DA and variational methods. Ensemble DA includes Monte Carlo based, ensemble Kalman methods (see, e.g., Evensen, 1994; Anderson, 2001; Tippett et al., 2003; Evensen et al., 2009), and “hybrid” methods that combine the Monte Carlo approach with optimization (see, e.g., Hamill and Snyder, 2000; Lorenc, 2003; Zhang et al., 2009; Buehner et al., 2013; Kuhl et al., 2013; Poterjoy and Zhang, 2015). Traditional variational DA (3D- or 4D-Var) uses optimization to find optimal state estimates, but does not randomly sample from a posterior distribution to form an ensemble (see, e.g., Talagrand and Courtier, 1987). Collectively, we will refer to ensemble and variational DA as “conventional DA,” because these methods have been deployed operationally for the past few decades.
@@
Recently, there has been a surge in interest in applying machine learning (ML) methods to geophysical modeling (see, e.g., Price et al., 2025; Lam et al., 2023; Kochkov et al., 2024), which raises the question of how ML tools might be used for DA. A first step might be to replace the physics-based forecast model within a conventional DA system with a data-driven ML emulator while continuing to employ conventional DA methods to merge forecasts with observations (see, e.g., Adrian et al., 2025). A more ambitious step might be to replace the entire conventional DA system with ML. This form of ML-based DA system can be implemented via diffusion models (a form of generative artificial intelligence (AI), see, e.g., Rozet and Louppe, 2023; Chung et al., 2023; Manshausen et al., 2024; Pathak et al., 2024; Qu et al., 2024; Li et al., 2025), deep neural networks (see, e.g., Allen et al., 2025) or by training a neural network that represents the functional of variational data assimilation, even in the absence of a training dataset (Keller and Potthast, 2024).
This paper explores the question: Can a diffusion model completely replace a conventional DA system? To answer this, we must first understand the subtle differences in the Bayesian posterior distributions that arise from different assumptions about the available prior distribution, which, in an ML/AI context, refers to the training set. An ensemble-based conventional DA system samples a Bayesian posterior distribution with a time-dependent prior that is informed by past observations. We will refer to this as a cycling prior because it propagates information from previous DA cycles to the current one. In contrast, we will demonstrate that a diffusion model that is trained on a long time-series of past weather data targets a different Bayesian posterior distribution with a constant, climatological prior. While some diffusion DA systems incorporate a “forecast” of some sort along with observations, we will show that this is still not the same as a cycling prior. Ultimately, we conclude that, except in rare cases (Bao et al. (2024)), ensemble DA and diffusion DA target different posterior distributions. We will formulate our arguments using a simple linear example with Gaussian errors, which allows for a clear and exact analysis without approximations. This leads to a crucial question: Which diffusion model formulation minimizes variance and forecast errors? Our analysis will unambiguously show that a diffusion model designed to sample the posterior distribution with a cycling prior, which is the one also targeted by current ensemble DA, is the optimal choice.
On the other hand, a DA method’s practical performance depends on how accurately it can sample the posterior distribution. In nonlinear systems, these distributions are often non-Gaussian, causing conventional DA methods that assume Gaussian distributions to produce significant errors (Morzfeld and Hodyss (2019)). This difficulty in sampling the posterior can obscure the theoretical differences between various DA methods. While important, analyzing these application-dependent sampling issues is beyond the scope of this paper. Instead, we focus on a more fundamental question: What specific variations of Bayesian posterior distributions are the various diffusion models sampling? By identifying this, we can clearly distinguish the algorithms in a way that is valid for all systems, whether they are linear, nonlinear, Gaussian, or non-Gaussian. The practical impact of non-Gaussianity is left for future work.
@@
The rest of this paper is organized as follows. In Section 2 we will introduce ensemble and variational methods as conventional DA systems. In Section 3 we introduce diffusion modeling and diffusion DA. In Section 4 we describe a linear, stochastic dynamical system that will allow us to clearly formulate all aspects of the DA problem analytically. We will apply different forms of diffusion DA systems to this dynamical model in order to reveal the prior, likelihood and posterior each system corresponds to. In Section 5 we provide a numerical illustration of the main results from Section 4. In Section 6 we provide a non-technical overview of the main results from sections 4 and 5. We close the manuscript with a summary of the major results and their conclusions in Section 7.
2 Ensemble and variational data assimilation
We briefly review fundamental concepts of ensemble and variational DA with an emphasis on how the methods target different Bayesian posterior distributions that are generated by different prior assumptions.
2.1 Ensemble data assimilation
@@
Ensemble DA is concerned with sampling a time-evolving Bayesian posterior distribution that describes the probability of a system state $\mathbf{x}_{k}$ at (discrete) time $k$, given observations $\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{k-1},\mathbf{y}_{k}$ up to time $k$:
@@
It is important to note here that information is propagated from cycle-to-cycle by the time evolving, or “cycling” prior $p(\mathbf{x}_{k}|\mathbf{y}_{1},\dots,\mathbf{y}_{k-1})$. In other words, the posterior of the previous cycle is used to generate the prior for the next cycle.
The ensemble Kalman filter (EnKF) generates ensembles that are approximately distributed according to the posterior distribution (1) as follows (see, e.g., Burgers et al., 1998; Evensen, 1994; Evensen et al., 2009). At cycle k1k-1, we have observations $\mathbf{y}_{1},\dots,\mathbf{y}_{k-1}$ and samples from $p(\mathbf{x}_{k-1}|\mathbf{y}_{1},\dots,\mathbf{y}_{k-1})$ in the form of an ensemble $\mathbf{x}_{k-1}^{(i)}$, where superscript $i=1,\dots,n_{e}$ indexes the ensemble members and where $n_{e}$ is the ensemble size. The EnKF makes a forecast for time $k$ by evolving the ensemble $\mathbf{x}_{k-1}^{(i)}$ forward in time using the model; the result is a forecast ensemble, viz.
@@
where $\mathcal{M}(\cdot)$ is a (possibly stochastic) forecast model, either physics-based or data-driven ML/AI. This forecast ensemble represents the cycling prior at time $k$, $p(\mathbf{x}_{k}|\mathbf{y}_{1},\dots,\mathbf{y}_{k-1})$, as a collection of samples. The forecast ensemble is updated to an analysis ensemble by employing stochastic ensemble generation (van Leeuwen (2020)), viz.
$\mathbf{x}_{ak}^{(i)}=\mathbf{x}_{fk}^{(i)}+\mathbf{K}(\mathbf{y}_{k}-(\mathbf{H}\mathbf{x}_{fk}^{(i)}+\bm{\varepsilon}^{(i)})),$ (3)
@@
$\mathbf{y}_{k}=\mathbf{H}\mathbf{x}_{k}+\bm{\varepsilon},$ (4)
where $\bm{\varepsilon}$ is a Gaussian random variable with mean $\mathbf{0}$ and covariance matrix $\mathbf{R}$; $\bm{\varepsilon}^{(i)}$ in (3) is the $i^{th}$ sample from the random variable $\bm{\varepsilon}$; the matrix
@@
is an ensemble approximation of the Kalman gain and $\mathbf{P}$ is the ensemble covariance, usually localized and inflated (see, e.g., Morzfeld and Hodyss, 2023; Whitaker and Hamill, 2012; Hodyss et al., 2016b; Gharamti et al., 2019)).
There also exist hybridized methods between the EnKF presented above and variational methods. Hybrid variational methods combine gradient-based optimization with an ensemble approach and target a similar Bayesian posterior distribution, but now randomly sample the state before the observation time (this is typically referred to as a “smoother”; see Hodyss et al. (2016a) for a review):
$p_{\text{EDA}}(\mathbf{x}_{k-1}|\mathbf{y}_{1},\dots,\mathbf{y}_{k})\propto p(\mathbf{x}_{k-1}|\mathbf{y}_{1},\dots,\mathbf{y}_{k-1})p(\mathbf{y}_{k}|\mathbf{x}_{k-1}),$ (6)
where $p(\mathbf{x}_{k-1}|\mathbf{y}_{1},\dots,\mathbf{y}_{k-1})$ is a cycling prior and $p(\mathbf{y}_{k}|\mathbf{x}_{k-1})$ is the likelihood, connecting the state at time k1k-1 to the observation at time $k$:
$\mathbf{y}_{k}=\mathbf{H}\mathcal{M}(\mathbf{x}_{k-1})+\bm{\varepsilon}.$ (7)
Specifically, an “ensemble of data assimilation’s” (EDA, Bonavita et al., 2012) maximizes randomized posterior distributions that are obtained by assuming that the cycling prior $p(\mathbf{x}_{k-1}|\mathbf{y}_{1},\dots,\mathbf{y}_{k-1})$ is Gaussian with mean $\bm{\mu}_{k-1}$ and flow-dependent covariance matrix $\mathbf{B}_{k-1}$, both generated from the ensemble at the previous cycle. The randomized negative log of the posterior distribution defines the loss function
$\mathcal{L}_{\text{EDA}}(\mathbf{x}_{k-1})=\frac{1}{2}\left\|\mathbf{R}^{-1/2}(\mathbf{y}_{k}-\mathbf{H}\mathcal{M}(\mathbf{x}_{k-1})+\bm{\varepsilon}^{(i)}\right\|^{2}+\frac{1}{2}\left\|\mathbf{B}_{k-1}^{-1/2}(\mathbf{x}_{k-1}-(\bm{\mu}_{k-1}+\bm{\eta}^{(i)}))\right\|^{2},$ (8)
where $\bm{\varepsilon}^{(i)}$ is Gaussian with mean zero and covariance matrix $\mathbf{R}$ and $\eta^{(i)}$ is the $i^{th}$ ensemble perturbation around the ensemble mean $\mu_{k-1}$. The minimum of the loss in equation (8) is taken as the $i^{th}$ posterior ensemble member, which in the linear, Gaussian case reduces precisely to the stochastic ensemble in equation (3). Thus, the EnKF and the EDA are quite closely related sampling strategies for generating ensembles with similar statistics.
For our purposes of interpreting diffusion DA in the context of Bayesian posterior distributions, the cycling nature of the priors used in the EnKF or EDA is important: The prior at the current cycle is directly informed by previous observations and the prior is changing from one cycle to the next as it is recalculated from the latest ensemble.
2.2 Variational data assimilation
Variational DA (3D- or 4D-Var) computes optimal state estimates by minimizing a loss function, but without the use of an ensemble. This point is worth emphasizing: classical 3D and 4D-var are not ensemble methods and hence do not randomly sample a posterior distribution. The statistical interpretation normally associated with variational methods is that they return only the mode of the posterior distribution (see Hodyss et al. (2016a)).
The loss function of 3D-Var is
$\mathcal{L}_{\text{3D-Var}}(\mathbf{x}_{k})=\frac{1}{2}\left\|\mathbf{R}^{-1/2}\left(\mathbf{y}_{k}-\mathbf{H}\mathbf{x}_{k}\right)\right\|^{2}+\frac{1}{2}\left\|\mathbf{B}^{-1/2}\left(\mathbf{x}_{k}-\bm{\mu}_{k}\right)\right\|^{2},$ (9)
where $\mathbf{R}$ and $\mathbf{B}$ are the observation error and background covariance matrices and where $\bm{\mu}_{k}$ is a background state (typically a forecast of some sort).
Similarly, 4D-Var is a smoother and as such uses the forward model $\mathcal{M}(\cdot)$ in (2) to estimate the state $\mathbf{x}_{k-1}$ before the observation time $k$ by minimizing the loss function
$\mathcal{L}_{\text{4D-Var}}(\mathbf{x}_{k-1})=\frac{1}{2}\left\|\mathbf{R}^{-1/2}\left(\mathbf{y}_{k}-\mathcal{H}(\mathcal{M}(\mathbf{x}_{k-1})\right)\right\|^{2}+\frac{1}{2}\left\|\mathbf{B}^{-1/2}\left(\mathbf{x}_{k-1}-\bm{\mu}_{k-1}\right)\right\|^{2}.$ (10)
Classical 3D and 4D-Var use time-independent background matrices that represent the forecast error in short-term forecasts of the length of the cycling interval. The celebrated NMC method (i.e. Parrish and Derber (1992)) was invented to make an estimate of precisely this quantity. For example, in the linear, Gaussian system we use below to illustrate our main results we will see that the prior error covariance matrix converges to a steady-state by balancing the tendency of observations to reduce the error with the tendency of the forecast model to increase the error. This steady-state prior covariance matrix is precisely the quantity that the NMC method is attempting to estimate using time-averaged statistics. In nonlinear/non-Gaussian systems, the prior covariance matrix does not reach a steady-state, but in weakly nonlinear/non-Gaussian systems, for which synoptic-scale weather is likely applicable, the prior error covariance matrix wanders around an equilibrium that could be interpreted as this steady-state. In this case, the NMC method attempts to find this time averaged (steady-state) covariance matrix.
3 Diffusion modeling and diffusion data assimilation
We briefly review the basics of diffusion modeling and how to use it for diffusion DA. We also briefly describe aspects of data-driven, ML/AI forecasting models and how these are related to our ideas.
@@ 3.1 Diffusion modeling
The goal in diffusion modeling is to construct a procedure to sample a random variable with probability density function (PDF), $p(\mathbf{x})$, given a sufficiently large number of samples from that distribution in the form of a training dataset (Sohl-Dickstein et al., 2015; Ho et al., 2020). The standard approach is to set up a forward process in the form of a simple stochastic differential equation (SDE) and to then reverse that process. In the forward process, we start with a sample from $p(\mathbf{x})$ and sequentially add noise to the sample. A neural network is trained on the samples (and the successively noisy versions of it) to enable a reverse process, which takes in noise and then outputs a sample from the PDF $p(\mathbf{x})$.
Following Karras et al. (2022), we use the stochastic differential equation (SDE)
@@
where $\beta_{t}$ is a standard Brownian motion (Wiener process), as the forward process. Its solution at time $t$ has the property that
$\mathbf{u}_{t}|\mathbf{u}_{0}\sim N(\mathbf{u}_{0},t^{2}\mathbf{I}),$ (12)
which means the the forward process amounts to adding Gaussian noise to samples drawn from $p(\mathbf{x})$. We emphasize that $t$ here is not a physical time, but a “virtual” time of the forward process.
Anderson (1982) constructs a reverse process that, when initialized with noise and integrated backwards in time from $t=T$ to $t=0$, generates samples from the PDF, $p(\bf{x})$, provided $T$ is large enough. The reverse process is governed by the SDE
@@
where $p_{t}(\mathbf{v})$ is the time evolution of the marginal PDF of (11) and $\overline{\beta}_{t}$ is a reverse-time Brownian motion. Thus, one can use the reverse process to generate samples from $p(\mathbf{x})$ by (i) drawing an initial condition from a Gaussian with mean zero and variance $T^{2}$ where $T$ is large; and (ii) simulating the reverse process (13) from time $T$ to time 0.
The term $\nabla\log p_{t}(\mathbf{v})$ is referred to as the “score function”, which can be computed via Tweedie’s formula (see Appendix A and Efron (2011))
@@
is a conditional expectation and, hence, the minimum variance estimator of $\mathbf{x}$ given $\mathbf{v}$. We can thus compute the conditional expectation as the minimizer of the loss function
[…]
@@
where $||\cdot||_{2}^{2}$ denotes the $\ell_{2}$ norm and where $D(\mathbf{v}_{t},t)$ is referred to as a “denoiser.” In practice, the denoiser is a neural network that is trained to predict $\mathbf{x}$ from $\mathbf{v}_{t}$ as follows. For each sample $\mathbf{x}$, we have noisy versions at various times in the forward process, which we can obtain by simply adding noise to the sample $\mathbf{x}$ according to the forward process. These “sample” and “noisy sample” pairs are used to train the denoiser. Once trained, the denoiser is used for simulating the reverse process
$\text{d}\mathbf{v}=-2\frac{D(\mathbf{v},t)-\mathbf{v}}{t}\text{d}t+\sqrt{2t}\text{d}\overline{\beta}_{t},$ (17)
initialized with noise, backwards in time to sample the PDF $p(\mathbf{x})$.
3.2 Diffusion data assimilation